OpenStudy (anonymous):

A straight line passes through points (a; 0) and (0; b), where a; b not equal to 0 . Prove that the line has the following equation: bx + ay = ab

OpenStudy (rulnick):

slope is -b/a (from the two points) y = (-b/a) x + b mult by a: ay = -bx + ab rearrange: bx + ay = ab Done.

OpenStudy (anonymous):

thanks!